OpenStudy (anonymous):

simplify sqrt 3c^2d (multiplied by) sqrt 6cd^2

OpenStudy (anonymous):

\[\sqrt{3c^{2}d}\times\sqrt{6cd^2}=\sqrt{18c^3d^3}=2\sqrt{3c^3d^3}\ for\ a,b,c,d \ being\ reals.\]